40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Jan 28 19:08:50 2009, Yoichi, Update, PSL, Laser is back (sort of) Improving.png
    Reply  Thu Jan 29 13:52:34 2009, Yoichi, Update, PSL, Laser is back (sort of) SlowDC.pdf
Message ID: 1256     Entry time: Wed Jan 28 19:08:50 2009     Reply to this: 1257
Author: Yoichi 
Type: Update 
Category: PSL 
Subject: Laser is back (sort of) 
Yoichi, Peter, Jenne

Summary:
We found that the chiller water is not going to the NPRO base. It was hot whereas it was cold when I touched it a few months ago.
I twisted the needle valve on the water line to the NPRO base. Then we heard gargling noise in the pipe and the water started to flow.
The laser power is now climbing up slowly. The noisiness of the MOPA output is reduced.

I will post more detailed entry explaining my theory of what actually happened later.
Attachment 1: Improving.png  155 kB  | Hide | Hide all
Improving.png
ELOG V3.1.3-